# ManiWAV: Learning Robot Manipulation from In-the-Wild Audio-Visual Data

### v1

@@ Abstract
Audio signals provide rich information for the robot interaction and object properties through contact. These information can surprisingly ease the learning of contact-rich robot manipulation skills, especially when the visual information alone is ambiguous or incomplete. However, the usage of audio data in robot manipulation has been constrained to teleoperated demonstrations collected by either attaching a microphone to the robot or object, which significantly limits its usage in robot learning pipelines. In this work, we introduce ManiWAV: an ‘ear-in-hand’ data collection device to collect in-the-wild human demonstrations with synchronous audio and visual feedback, and a corresponding policy interface to learn robot manipulation policy directly from the demonstrations. We demonstrate the capabilities of our system through four contact-rich manipulation tasks that require either passively sensing the contact events and modes, or actively sensing the object surface materials and states. In addition, we show that our system can generalize to unseen in-the-wild environments, by learning from diverse in-the-wild human demonstrations.
@@
Fig 1: Contact sound reveals rich information. More specifically, (a) contact events such as eraser touching the whiteboard; (b) different contact modes such as spatula poking on the side of the bagel versus sliding below the bagel; (c) surface material such as the furry (‘loop’) and rough (‘hook’) side of velcro tapes; (d) object states such as whether the cup contains objects or not.
@@ 1 Introduction
  • Surface materials: Audio signals can be used to characterize the surface material through contact with the object. In contrast, either image sensors or vision-based tactile sensors require high spatial resolution to capture the subtle texture difference (e.g. the ‘hook’ and ‘loop’ side of velcro tapes) (Fig. 1 c).
@@
  • Object states and properties: With indirect contacts, audio signals can provide complementary information about the object state and physical properties beyond visual observation (Fig. 1 d).
@@
Second, audio data is scalable for data collection and policy learning. This is because acoustic sensors (i.e. contact microphones) are cheap, robust, and readily available to purchase. Audio signals also have standardized coding formats that can be easily integrated into existing video recording and storage pipeline (e.g. MP4 files). These nice properties make it possible to collect audio in-the-wild with low-cost data collection devices such as a hand-held gripper, without the need of a robot. On the other hand, alternative ways to sense contact such as tactile sensors are relatively more expensive, fragile, and requires expert knowledge to use.
Given the richness and scalability of audio data, we propose a versatile robot learning system, ManiWAV, that leverages audio feedback for contact-rich robot manipulation tasks:
  • On the hardware side, we propose a portable hand-held device for in-the-wild data collection. As shown in Fig. 2, we introduce an “ear-in-hand” design that allows human to easily demonstrate a wide range of manipulation tasks with synchronous vision and audio feedback. The hand-held gripper also naturally provides haptic feedback (e.g. when wiping on whiteboard) during data collection, which is hard to obtain with teleoperation. To capture contact audio siganls with a higher signal-to-noise ratio, we choose to use a piezoelectric contact microphone with textured high-grip surface attached on top.
@@
  • On the algorithm side, one key challenge is to bridge the audio domain gap between in-the-wild data and actual robot deployment due to test-time noises (Fig. 2 b). To achieve this goal, we propose a data augmentation strategy that encourages learning of task-relevant audio representation. In addition, we propose an end-to-end sensorimotor learning network to encode and fuse the vision and audio information, with diffusion head for action prediction. We show that this network design helps learn better audio-visual representation from multimodal human demonstrations.
@@
We demonstrate the capability of our proposed system on four contact-rich manipulation tasks: wipe shape from whiteboard, flip bagel with spatula, pour objects from cup, and strap wires with velcro tape. We also show that our system is able to generalize to unseen in-the-wild environments, by leveraging in-the-wild data collected from diverse environments.
@@
Fig 2: Ear-in-hand gripper for in-the-wild data collection. (a) The handheld design naturally provide haptic feedback to the demonstrator during contact-rich tasks (e.g. wiping), which is otherwise hard to obtain via teleoperation. Contact microphone captures high frequency audio feedback that is recorded simultaneously with images. High friction tape is applied on top to augment the signals. (b) shows the domain gap between training and deployment data. (c) shows policy learned from in-the-wild data directly deployed on the robot.
@@ 2 Related Work
Tactile Sensing for Contact-rich Manipulation. As a functional equivalent to the human touch modality, tactile sensing has long been studied to provide feedback for the robot’s physical interactions [1]. Ranging from six-axis force/torque sensors [2, 3] to camera-based tactile sensors [4, 5, 6, 7, 8, 9] and tactile skins [10, 11, 12, 13], tactile sensors take various forms and provide information of contact forces, object geometry, etc. Many recent works have incorporated tactile feedback into robot learning pipeline for contact-rich manipulation tasks [14, 15, 16, 17, 18, 19] by learning better visuo-tactile representations. However, most tactile sensors are limited in their reproducibility given the cost and requirement for domain knowledge to use. In this work, we find that acoustic sensors such as piezoelectric contact microphone can provide a wide range of alternative tactile feedback, at a much lower cost and availability.
Acoustic Sensing in Controlled Environments. Sound is an important information carrier of the physical environment. Acoustic sensing can be categorized into active and passive. Active acoustic sensing is done by emitting a vibration waveform with a speaker, which is then received by a microphone. Prior works have embedded active acoustic sensors on objects [20], robot arms [21, 22], parallel gripper [23] or soft finger [24] to sense object material, shape, and contact interactions. Passive acoustic sensing captures sounds generated from interactions, and prior works have shown that including audio as input to end-to-end robot learning algorithms improves performance of several manipulation tasks [25, 26, 27, 28]. However, the collection of audio data in prior works requires a controlled environment, where the data is collected through teleoperation, with microphone attached to either the robot or the object. To address this limitation, we propose an ‘ear-in-hand’ gripper design with passive acoustic sensing to collect human demonstrations without the need of a robot, making it promising to collect in-the-wild audio-visual data at lower cost of time.
Policy Learning from Multisensory Data. Multisensory observations (e.g, vision, audio, tactile) allows the robot to better perceive the physical environment and guide action planning [14, 29, 30, 31]. However, many prior works require pre-defined state abstractions [32, 33] to learn the control policy and are thus limited in their ability to generalize to new tasks. Recently end-to-end models have been proposed to take in multisensory inputs and output robot actions through a behavior cloning approach [28, 26, 27, 25]. We extend upon prior works and propose an audio data augmentation strategy to bridge the domain gap between in-the-wild data and on-robot data during deployment, and an end-to-end policy learning network to effectively learn task-relevant audio-visual representations from multimodal human demonstrations.
@@ 3 Method
We propose a data collection and policy learning framework to learn contact-rich manipulation tasks from vision and audio. On the data collection front, our goal is to easily collect in-the-wild demonstrations with clean and salient contact signals. To achieve this end, we propose an “ear-in-hand” data collection device to collect synchronized vision and audio data with a GoPro camera and a contact microphone.
@@ 3.1 Ear-in-Hand Hardware Design
Our data collection device is built on top of Universal Manipulation Interface (UMI) [34]. UMI is a portable and low-cost hand-held gripper designed to collect rich human demonstrations in the wild. The collected data can be used to train a visuomotor policy directly deployable on a robot.
We redesign the 3D printed parallel jaw gripper on the device to embed a piezoelectric contact microphone under high-friction grip tape wrapped around the finger. The microphone is connected to the 3.5mm external mic port on the GoPro camera media mod. Fig. 2 (a) shows the hand-held gripper design. Audio is recorded at 48000 Hz and stored with 60Hz image data synchronously as MP4 files. During robot deployment, the same parallel jaw gripper with embedded microphone is mounted on a UR5 robot arm, shown in Fig. 2 (c). The images and audio are streamed real time through an Elgato HD60 X external capture card into a Ubuntu 22.04.3 desktop.
@@ 3.2 Policy Design
We propose an end-to-end closed-loop sensorimotor learning model that takes in RGB images and audio, and output 10-DoF robot actions. (end effector positions, end effector orientation represented in 6D [35], and 1D gripper openness).
Audio Data Augmentation. One key challenge is that the audio signals received during real time robot deployment is very different from the data collected by the hand-held gripper, resulting in a large domain gap between our training and test scenarios, as illustrated in Fig. 2 (b). This is mostly because 1) nonlinear robot motor noises during deployment, 2) out-of-distribution sounds generated by the robot interaction. (e.g. accidentally collide with an object).
@@
To address the domain gap, the key is to augment the training data with noises and guide the model to focus on the invariant task-relevant signals and ignore unpredictable noises. In particular, we randomly sample audio as background noises from ESC-50 [36]. The sounds are normalized to the same scale as the collected sound in the training dataset. We also record 10 samples of robot motor noises under randomly sampled trajectories with same contact microphone location as deployment time. The background noises and robot noises are overlayed to the original audio signal, each with a probability of 0.5. In our experiments, we show that this simple yet effective approach yields better policy performance by enforcing the inductive bias of the model on task-relevant audio signals.
Fig 4: Attention Visualization. Interestingly, we find that a policy co-trained with audio attends more on the task-relevant regions (shape of drawing or free space inside the pan). In contrast, the vision only policy often overfits to background structures as an shortcut to estimate contact (e.g., the edge of the whiteboard, table, and room structures).
@@
Audio Encoder. We use the audio spectrogram transformer (AST) [38] to encode the audio input. AST, similar to a ViT model, leverages attention mechanism to learn better audio representation from spectrogram patches. The intuition behind using a transformer encoder instead of a CNN-based encoder as seen in prior works [26, 28, 25] is that the ‘shift invariance’ that CNN leverages is less suitable to audio spectrograms, as shift in either the time and frequency domain can significantly change the information. In our experiment, we show that training the transformer encoder from scratch outperforms both pre-trained and from-scratch CNN models.
The audio singal (from last 2-3 seconds depending on the task) is first resampled from 48kHz to 16kHz and converted to a log mel spectrogram using FFT size and window length of 400, hop length of 160, and 64 mel filterbanks. The log-mel spectrograms are linearly normalized to range [1,1]11[-1,1][ - 1 , 1 ]. We use the classification token feature extracted from the last hidden layer of the AST encoder.
Sensory Fusion. We fuse the vision and audio features using a transformer encoder in a similar fashion as Li et al. [28] to leverage the attention mechanism to weigh the features adaptively at different stages of the task (e.g. vision is important for moving to the target object, whereas audio is important during contact). We concatenate the output features and downsample the dimension to 768 with a linear projection layer. Finally we concatenate the end effector poses (20 Hz) from the past 2 timesteps to the audio-visual feature.
Policy Learning. To model the multimodality intrinsic to human demonstrations, we choose to use a diffusion model with UNet encoders as proposed by Chi et al. [39] as the policy head, conditioned on the observation representation mentioned above in each denoising step. The entire model (Fig. 3), including the above mentioned encoders, is end-to-end trained using the noise prediction MSE loss on future robot trajectories of 16 steps.
Audio Latency Matching. During data collection, the vision and audio data are synchronized during recording through the GoPro. During deployment, we calibrated the audio latency to be 0.23; details can be found in appendix. We adopt an approach similar to Chi et al. [34] to compensate for this delay.
@@ 4 Evaluation
We study four contact-rich manipulation tasks to show the different capabilities of audio feedback, such as detecting contact events and modes (flipping and wiping), sense object state (pouring) and surface material (taping). In each task, we test the policy under different scenarios and compare with alternative approaches to validate the robustness and generalizability of our approach.
Fig 5: Flipping Evaluation. Up: Different test scenarios including both in-lab and in-the-wild environments. Bottom: Typical failure cases and task success rate. More results can be found on the project website. The details (e.g. failure case) for each rollout can be found in appendix.
@@ 4.1 Flipping Task
The robot is tasked to flip a bagel in a pan from facing down to facing upward using a spatula. To perform this task successfully, the robot needs to sense and switch between different contact modes – precisely insert the spatula between the bagel and the pan, maintain the contact while sliding, and start to tilt up the spatula when the bagel is in contact with the edge of the pan.
Data collection: We collect two types of demonstrations for this task: 115 in-the-lab demonstrations, and an additional 274 in-the-wild demonstrations collected in 6 different environments using different pans. The hand-held gripper naturally provide haptic feedback upon contact which makes the collection more intuitive comparing to teleoperation. Each demonstration takes about 4-6 seconds to collect.
Test Scenarios: We run 20 rollouts for each policy. To ensure fair comparison, we use the same set of robot and object configurations for evaluations between different methods. We achieve the same object configuration by overlaying their position with respect to a captured image in the camera view. The test configurations can be grouped into four categories:
  • T1: Variations in task configuration: different initial robot and object configurations (14 / 40).
  • T2: Audio perturbation by playing noises in the background (2 / 40).
  • T3: Generalization to unseen table height (4 / 40).
  • T4: Generalization to two unseen in-the-wild environments: a black stovetop and a white countertop, the later is more challenging due to unstructured background and lack of similar training data (20 / 40).
@@
  • MLP policy: using a MLP with three hidden layers (following Li et al. [28]) instead of action diffusion. The model takes the observation representation and outputs the future action trajectory.
@@
  • In-the-wild: For in-the-wild evaluation (T4), we compare our model trained with in-the-wild demonstrations, only in-lab demonstrations, and a [Vision only] baseline trained on in-the-wild data. Details on the in-the-wild dataset can be found in appendix.
Findings: The quantitative result and typical failure cases are visualized in Fig. 5. Our key findings are: 1) Action diffusion yields better behavior comparing to MLP, yet the [MLP policy] that takes in audio data still significantly outperforms [Vision only] diffusion policy. Action diffusion better captures the multimodality in human demonstrations, for example, one might approach the bagel from different directions in different demonstrations. 2) Transformer audio encoder outperforms a CNN-based encoder. We find that training a transformer encoder from scratch yields good performance comparing to using a CNN-based encoder, as shown in the [ResNet] and [AVID] baseline. This is likely because the self-attention mechanism in transformer allows the model to focus more on the frequency regions of task-relevant signals in the spectrogram. 3) In-the-wild data enables generalization to unseen in-the-wild environments. As shown in Fig. 5, policy trained on in-the-wild data significantly outperforms in-lab data on two unseen environments, as the scene diversity in in-the-wild data allows the policy to generalize better to new environments.
Fig 6: Wiping Evaluation. Up: Different test scenarios. Bottom: Typical failure cases and task success rate. [Vision only] policy often fails to maintain proper contact (e.g., either press too hard into the broad or float). [MLP fusion] policy often fails to fully wipe out the drawing and terminate early.

### v2

@@ Abstract
Audio signals provide rich information for the robot interaction and object properties through contact. This information can surprisingly ease the learning of contact-rich robot manipulation skills, especially when the visual information alone is ambiguous or incomplete. However, the usage of audio data in robot manipulation has been constrained to teleoperated demonstrations collected by either attaching a microphone to the robot or object, which significantly limits its usage in robot learning pipelines. In this work, we introduce ManiWAV: an ‘ear-in-hand’ data collection device to collect in-the-wild human demonstrations with synchronous audio and visual feedback, and a corresponding policy interface to learn robot manipulation policy directly from the demonstrations. We demonstrate the capabilities of our system through four contact-rich manipulation tasks that require either passively sensing the contact events and modes, or actively sensing the object surface materials and states. In addition, we show that our system can generalize to unseen in-the-wild environments by learning from diverse in-the-wild human demonstrations.
@@
Fig 1: Contact sound reveals rich information. More specifically, (a) contact events, such as eraser touching the whiteboard; (b) different contact modes, such as spatula poking on the side of the bagel versus sliding below the bagel; (c) surface material, such as the furry (‘loop’) and rough (‘hook’) side of velcro tapes; (d) object states, such as whether the cup contains objects or not.
@@ 1 Introduction
  • Surface materials: Audio signals can be used to characterize the surface material through contact with the object. In contrast, either image sensors or vision-based tactile sensors require high spatial resolution to capture the subtle texture difference (e.g., the ‘hook’ and ‘loop’ side of velcro tapes) (Fig. 1 c).
@@
  • Object states and properties: Without the need for direct contact, audio signals can provide complementary information about the object state and physical properties beyond visual observation (Fig. 1 d).
@@
Second, audio data is scalable for data collection and policy learning. This is because acoustic sensors (i.e., contact microphones) are cheap, robust, and readily available to purchase. Audio signals also have standardized coding formats that can be easily integrated into existing video recording and storage pipelines (e.g., MP4 files). These nice properties make it possible to collect audio in the wild with low-cost data collection devices, such as a hand-held gripper, without the need for a robot. On the other hand, alternative ways to sense contact, such as tactile sensors, are relatively more expensive, fragile, and require expert knowledge to use.
Given the richness and scalability of audio data, we propose a versatile robot learning system, ManiWAV, that leverages audio feedback for contact-rich robot manipulation tasks. Building upon the portable hand-held data collection device UMI [1], we redesign one gripper finger to embed a piezoelectric contact microphone that senses audio vibrations through contact with solid objects. The audio signals can be easily streamed to the GoPro camera through a mic port and stored synchronously with vision data in MP4 files. With this intuitive design, one can demonstrate a wide range of manipulation tasks with synchronous vision and audio feedback at a low time and maintenance cost.
@@
To learn from the collected demonstrations, one key challenge is to bridge the audio domain gap between in-the-wild data and actual robot deployment due to test-time noises (Fig. 2 b). To achieve this goal, we propose a data augmentation strategy that encourages learning of task-relevant audio representation. In addition, we propose an end-to-end sensorimotor learning network to encode and fuse the vision and audio data, with a diffusion policy [2] head for action prediction.
@@
We demonstrate the capability of our proposed system on four contact-rich manipulation tasks: wipe shape from whiteboard, flip bagel with spatula, pour objects from cup, and strap wires with velcro tape. We also show that our system can generalize to unseen in-the-wild environments by leveraging in-the-wild data collected from diverse environments.
@@
Fig 2: Ear-in-hand gripper for in-the-wild data collection. (a) The handheld design naturally provides haptic feedback to the demonstrator during contact-rich tasks (e.g., wiping), which is otherwise hard to obtain via teleoperation. Contact microphone captures high-frequency audio feedback that is recorded simultaneously with images. High friction tape is applied on top to augment the signals. (b) shows the domain gap between training and deployment data. (c) shows policy learned from in-the-wild data directly deployed on the robot.
@@ 2 Related Work
Tactile Sensing for Contact-rich Manipulation. As a functional equivalent to the human touch modality, tactile sensing has long been studied to provide feedback for the robot’s physical interactions [3]. Ranging from six-axis force/torque sensors [4, 5] to camera-based tactile sensors [6, 7, 8, 9, 10, 11] and tactile skins [12, 13, 14, 15], tactile sensors take various forms and provide information of contact forces [16], object geometry [17, 18], etc. Many recent works have incorporated tactile feedback into robot learning pipelines for contact-rich manipulation tasks by learning better visuo-tactile representations [19, 20, 21, 22, 23, 24]. However, most tactile sensors are limited in their reproducibility given the cost and requirement for domain knowledge to use. In this work, we find that acoustic sensors such as piezoelectric contact microphones can provide alternative tactile feedback at a much lower cost and higher availability.
Acoustic Sensing in Controlled Environments. Sound is an important information carrier of the physical environment. Acoustic sensing can be categorized into active and passive. Active acoustic sensing is done by emitting a vibration waveform with a speaker, which is then received by a microphone. Prior works have embedded active acoustic sensors on objects [25], robot arms [26, 27], parallel gripper [28] or soft finger [29] to sense object material, shape, and contact interactions. Passive acoustic sensing captures sounds generated from interactions, and prior works have shown that including audio as input to end-to-end robot learning algorithms improves the performance of manipulation tasks [30, 31, 32, 33]. However, the collection of audio data in prior works requires a controlled environment, where the data is collected through teleoperation, with sensors attached to either the robot or the object. To address this limitation, we propose an ‘ear-in-hand’ gripper design with passive acoustic sensing to collect human demonstrations without the need for a robot, making it promising to collect in-the-wild audio-visual data at a lower cost of time.
Policy Learning from Multisensory Data. Multisensory observations (e.g., vision, audio, tactile) allow robots to better perceive the physical environment and guide action planning [19, 34, 35, 36]. However, many prior works require pre-defined state abstractions [37, 38] to learn the control policy and are thus limited in their ability to generalize to new tasks. Recently, end-to-end models have been proposed to take in multisensory inputs and output robot actions through a behavior cloning approach [33, 31, 32, 30]. We extend upon prior works and propose an audio data augmentation strategy to bridge the domain gap between in-the-wild data and robot data during deployment, together with an end-to-end policy learning network to effectively learn task-relevant audio-visual representations from multimodal human demonstrations.
@@ 3 Method
We propose a data collection and policy learning framework to learn contact-rich manipulation tasks from vision and audio. On the data collection front, our goal is to easily collect in-the-wild demonstrations with clean and salient contact signals. To achieve this, we extend the UMI [1] data collection device by embedding a piezoelectric contact microphone in one gripper finger to stream audio data synchronously with the vision data through the GoPro camera mic port.
@@ 3.1 Ear-in-Hand Hardware Design
Our data collection device is built on the Universal Manipulation Interface (UMI) [1]. UMI is a portable and low-cost hand-held gripper designed to collect human demonstrations in the wild. The collected data can be used to train a visuomotor policy that is directly deployable on a robot.
We redesign the 3D-printed parallel jaw gripper on the device to embed a piezoelectric contact microphone under high-friction grip tape wrapped around the finger. The microphone is connected to the 3.5mm external mic port on the GoPro camera media mod. Fig. 2 (a) shows the hand-held gripper design. Audio is recorded at 48000 Hz and stored with 60Hz image data synchronously as MP4 files. During robot deployment, the same parallel jaw gripper with embedded contact microphone is mounted on a UR5 robot arm, shown in Fig. 2 (c). The images and audio are streamed in real-time through an Elgato HD60 X external capture card into a Ubuntu 22.04.3 desktop.
@@ 3.2 Policy Design
We propose an end-to-end closed-loop sensorimotor learning model that takes in RGB images and audio, and outputs 10-DoF robot actions (including end effector positions, end effector orientation represented in 6D [39], and 1D gripper openness).
Audio Data Augmentation. One key challenge is that the audio signals received during real-time robot deployment are very different from the data collected by the hand-held gripper, resulting in a large domain gap between the training and test scenarios, as illustrated in Fig. 2 (b). This is mostly because of 1) nonlinear robot motor noises during deployment and 2) out-of-distribution sounds generated by the robot interaction. (e.g., accidentally colliding with an object).
@@
To address the domain gap, the key is to augment the training data with noises and guide the model to focus on the invariant task-relevant signals and ignore unpredictable noises. In particular, we randomly sample audio as background noises from ESC-50 [40]. The sounds are normalized to the same scale as the collected sound in the training dataset. We also record 10 samples of robot motor noises under randomly sampled trajectories with the same contact microphone location as deployment time. The background noises and robot noises are overlayed to the original audio signal, each with a probability of 0.5. In our experiments, we show that this simple yet effective approach yields better policy performance by enforcing the inductive bias of the model on task-relevant audio signals.
Fig 4: Attention Visualization. Interestingly, we find that a policy co-trained with audio attends more to the task-relevant regions (shape of drawing or free space inside the pan). In contrast, the vision-only policy often overfits to background structures as a shortcut to estimate contact (e.g., the edge of the whiteboard, table, and room structures).
@@
Audio Encoder. We use the audio spectrogram transformer (AST) [42] to encode the audio input. AST, similar to a ViT model, leverages the attention mechanism to learn better audio representation from spectrogram patches. The intuition behind using a transformer encoder instead of a CNN-based encoder as seen in prior works [31, 33, 30] is that the ‘shift invariance’ that CNN leverages is less suitable to audio spectrograms, as shifts in either the time and frequency domain can significantly change the audio information. In our experiment, we show that training the transformer encoder from scratch outperforms both pre-trained and from-scratch CNN models.
The audio signal (from the last 2-3 seconds, depending on the task) is first resampled from 48kHz to 16kHz, which helps filter high-frequency noises and increase the frequency resolution of task-relevant signals on the spectrogram. The waveform is then converted to a log mel spectrogram using FFT size and window length of 400, hop length of 160, and 64 mel filterbanks. The log-mel spectrograms are linearly normalized to range [1,1]11[-1,1][ - 1 , 1 ]. We use the classification token feature extracted from the last hidden layer of the AST encoder.
Sensory Fusion. We fuse the vision and audio features using a transformer encoder in a similar fashion as Li et al. [33] to leverage the attention mechanism to weigh the features adaptively at different stages of the task (e.g., vision is important for moving to the target object, whereas audio is important upon contact). We concatenate the output features and downsample the dimension to 768 with a linear projection layer. Finally, we concatenate the end effector poses (20 Hz) from the past 2 timesteps to the audio-visual feature.
Policy Learning. To model the multimodality intrinsic to human demonstrations, we choose to use a diffusion model with UNet encoders as proposed by Chi et al. [2] as the policy head, conditioned on the observation representation mentioned above in each denoising step. The entire model (Fig. 3), including the above-mentioned encoders, is end-to-end trained using the noise prediction MSE loss on future robot trajectories of 16 steps.
Audio Latency Matching. During data collection, the vision and audio data are synchronized when recording through GoPro. During deployment, we calibrated the audio latency to be 0.23; details can be found in the appendix. We adopt an approach similar to Chi et al. [1] to compensate for this delay.
@@ 4 Evaluation
We study four contact-rich manipulation tasks to show the different advantages of learning from audio feedback, such as detecting contact events and modes (flipping and wiping), or sensing object state (pouring) and surface material (taping). In each task, we test the policy under different scenarios and compare with alternative approaches to validate the robustness and generalizability of our approach.
Fig 5: Flipping Evaluation. Up: On the left, we show the in-lab test scenarios. We train the policy with in-lab demonstrations collected in the same environment as inference time. We study three types of test-time variations: including different robot and object configurations (T1), audio perturbation when we play different types of noises in the environment (T2), and different table heights than the one in training data (T3). On the right, we show the two unseen environments for the in-the-wild generalization test. Bottom: Typical failure cases and task success rate. The details (e.g., failure cases) for each rollout can be found in the appendix.
@@ 4.1 Flipping Task
The robot is tasked to flip a bagel in a pan from facing down to facing upward using a spatula. To perform this task successfully, the robot needs to sense and switch between different contact modes – precisely insert the spatula between the bagel and the pan, maintain contact while sliding, and start to tilt up the spatula when the bagel is in contact with the edge of the pan.
Data collection: We collect two types of demonstrations for this task: 283 in-lab demonstrations and an additional 274 in-the-wild demonstrations collected in 7 different environments using different pans.
Test Scenarios: We run 20 rollouts for each policy. To ensure a fair comparison, we use the same set of robot and object configurations for evaluations between different methods. We achieve the same object configuration by overlaying their position with respect to a captured image in the camera view. The test configurations can be grouped into four categories.
  • T1: Variations in task configuration: different initial robot and object configurations (14 / 20).
  • T2: Audio perturbation by playing different types of noises in the background (2 / 20).
  • T3: Generalization to unseen table height (4 / 20).
  • T4: Generalization to two unseen in-the-wild environments: a black stovetop and a white countertop, the latter is more challenging due to unstructured background and lack of similar training data.
@@
  • MLP policy: using an MLP with three hidden layers (following Li et al. [33]) instead of action diffusion. The model takes the observation representation and outputs the future action trajectory.
@@
  • In-the-wild: For in-the-wild evaluation (T4), we compare our model trained with in-the-wild demonstrations (blue bar in Fig. 5 bottom right chart), only in-lab demonstrations (red bar in Fig. 5 bottom right chart), and a [Vision only] baseline trained on in-the-wild data (yellow bar in Fig. 5 bottom right chart). Details on the in-the-wild dataset can be found in the appendix.
Findings: The quantitative result and typical failure cases are visualized in Fig. 5. Our key findings are: 1) Action diffusion yields better behavior compared to MLP, yet the [MLP policy] that takes in audio data still significantly outperforms [Vision only] diffusion policy. Action diffusion better captures the multimodality in human demonstrations; for example, one might approach the bagel from different directions in different demonstrations. 2) Transformer audio encoder outperforms a CNN-based encoder. We find that training a transformer encoder from scratch yields good performance compared to using a CNN-based encoder, as shown in the [ResNet] and [AVID] baseline. This is likely because the self-attention mechanism in the transformer allows the model to focus more on the frequency regions of task-relevant signals in the spectrogram. 3) In-the-wild data enables generalization to unseen in-the-wild environments. As shown in Fig. 5, policy trained on in-the-wild data significantly outperforms in-lab data on two unseen environments, as the scene diversity in in-the-wild data allows the policy to generalize better to new environments.
Fig 6: Wiping Evaluation. Up: Different test scenarios. Bottom: Typical failure cases and task success rate. [Vision only] policy often fails to maintain proper contact (e.g., either press too hard into the broad or float). [MLP fusion] policy often fails to fully wipe out the drawing and terminates early.